### v2

@@ I Introduction
For simplicity we refer to this as “measuring $H$.” Without loss of generality we take $H$ to be an $n$-qubit Hamiltonian written as a linear combination of Pauli strings
@@
Figure 1: Measuring an $n=40$ qubit spinless, two-dimensional Fermi-Hubbard Hamiltonian (5) by grouping terms into $k$-commuting sets for $1\leq k\leq n$. The left axis (blue) shows the number of groups as a function of $k$ and the right axis (green) shows $\hat{R}$ defined in (6) as a function of $k$. As can be seen, the number of groups generally decreases as $k$ increases, and $\hat{R}$ increases as $k$ increases. Additionally, the plot shows a value $k^{*}<n$ at which $\hat{R}$ saturates. The inset shows this value $k^{*}(n)$ as a function of system size for $n=40,80,$ and $120$ qubit Fermi-Hubbard Hamiltonians for which we see linear scaling.
Motivated by the desire to have short-depth circuits due to hardware constraints, the notion of qubit-wise commutativity was developed [qubit-wise-commuting] (and even prior to this reference under the name of “tensor product basis sets” [Kandala_Mezzacapo_Temme_Takita_Brink_Chow_Gambetta_2017], although the term qubit-wise commutativity is most commonly used). Two Pauli strings $P$ and $Q$ are said to qubit-wise commute if $[P_{i},Q_{i}]=0$ for all $i\in[n]$. If $P$ and $Q$ qubit-wise commute, they can be measured together in a depth-one quantum circuit. The catch is that, given a generic Hamiltonian (2), not all Pauli strings will qubit-wise commute. Approximate optimization algorithms are employed to group Paulis into qubit-wise commuting sets such that each set can be measured with a single depth-one circuit. This provides a trade off, decreasing circuit depth at the cost of increasing the number of circuits needed to be run.
Thus there are two notions of commutativity used in measuring $n$-qubit Hamiltonians: qubit-wise commuting which results in $N\geq 1$ circuits of additional depth $d=1$, and fully commuting which results in $N=1$ circuit of additional depth $d=O(n^{2}/\log n)$. In this paper we introduce a new notion of commutativity which interpolates between these two. The new notion is simply to consider commutativity on blocks of size $k\in\mathbb{N}$, and we call this $k$-commutativity. For example, we say Paulis $XXYY$ and $ZZXX$ two-commute since $[XX,ZZ]=0$ (the first two-qubit block) and $[YY,XX]=0$ (the second two-qubit block). Qubit-wise commutativity is recovered by $k=1$ and (full) commutativity is recovered by $k=n$. The advantage, with respect to measuring Hamiltonians, is the potential to reduce measurement complexity by allowing for increased circuit depth in a smooth manner. An example of this is shown in Fig. 1 in which $k$-commutativity is applied to measuring the two-dimensional Fermi-Hubbard Hamiltonian. As $k$ increases, the number of groups decreases, and the value of $\hat{R}$ (a ratio of the advantage of grouping compared to not grouping, to be defined later) increases, demonstrating reduced measurement complexity.
@@ II $k$-commutativity
Let $P=P_{0}\otimes\cdots\otimes P_{n-1}$ be an $n$-qubit Pauli string. Given integers $a<b$, let $P_{a:b}$ denote $P$ restricted to $a$ through $b-1$, i.e. $P_{a}\otimes\cdots\otimes P_{b-1}$. For convenience we consider $P_{a},...,P_{-1}$ to be identity if $a<0$, and similarly $P_{n},...,P_{b-1}$ to be identity if $b>n-1$. We define $k$-commutativity as follows.
@@
Two $n$-qubit Pauli strings $P=P_{0}\otimes\cdots\otimes P_{n-1}$ and $Q=Q_{0}\otimes\cdots\otimes Q_{n-1}$ are said to $k$-commute for integer $1\leq k\leq n$ if and only if $[P_{ik:(i+1)k},Q_{ik:(i+1)k}]=0$ for all $i=0,...,\lfloor n/k\rfloor-1$. If $P$ and $Q$ $k$-commute, we write $[P,Q]_{k}=0$.
Note that this definition assumes an ordering of the operators and throughout the paper we consider the ordering to be specified when evaluating $k$-commutativity.
As mentioned, we note that $k=1$ corresponds to qubit-wise commutativity [qubit-wise-commuting] and $k=n$ corresponds to full commutativity. Due to its definition, $k$-commutativity inherits the usual properties of commutativity. In particular, $k$-commutativity is reflexive ($[P,Q]_{k}=0$ implies $[Q,P]_{k}=0$) but not transitive (if $[P,Q]_{k}=0$ and $[Q,R]_{k}=0$, then $[P,R]_{k}$ could be zero or nonzero).
@@
Our next properties concern the number of Pauli strings that $k$-commute with a given Pauli. For this we show a result for a slightly more general notion of $k$-commutativity in which the block sizes can vary. In particular, given two $n$-qubit Pauli strings $P$ and $Q$, and $k_{1},...,k_{m}\in\mathbb{N}$ such that $k_{1}+\cdots+k_{m}=n$, we say that $P$ and $Q$ $(k_{1},...,k_{m})$-commute if and only if $[P_{a_{j}:b_{j}},Q_{a_{j}:b_{j}}]=0$ for every $j=1,\dots,m$, where $b_{j}=\sum_{\ell=1}^{j}k_{j}$ and $a_{j}=b_{j}-k_{j}$.
@@
For every $j=1,\dots,m$, define $b_{j}=\sum_{\ell=1}^{j}k_{j}$, and $a_{j}=b_{j}-k_{j}$. If $Q$ is a $n$-qubit Pauli that $(k_{1},\dots,k_{m})$-commutes with $P$, then for every $j$, we have $[P_{a_{j}:b_{j}},Q_{a_{j}:b_{j}}]=0$. By [sarkar2021sets, Lemma 3], the number of distinct choices (modulo phase factors) for $Q_{a_{j}:b_{j}}$ is $24^{k_{j}}/$ 2. This gives that the total number of distinct possible choices (modulo phase factors) for $Q$ is $\prod_{j=1}^{m}(4^{k_{j}}/2)=4^{n}/2^{m}$. ∎
@@
Last, we note that the procedure for constructing a Clifford circuit to measure a set of $k$-commuting Paulis is identical to the case for (full) commutation [Crawford_Straaten_Wang_Parks_Campbell_Brierley_2021, Gokhale_Angiuli_Ding_Gui_Tomesh_Suchara_Martonosi_Chong_2020], just with the diagonalization procedure performed in $\lfloor n/k\rfloor$ blocks of size $k$ instead of one block of size $n$.
@@ III Measurement reduction
As an initial test we consider the spinless Fermi-Hubbard Hamiltonian
$H=-t\sum_{\langle i,j\rangle}(a^{\dagger}_{i}a_{j}+a^{\dagger}_{j}a_{i})+g\sum_{\langle i,j\rangle}a^{\dagger}_{i}a_{i}a^{\dagger}_{j}a_{j}-\mu\sum_{i}a_{i}^{\dagger}a_{i}$ (5)
on a two-dimensional grid. Here, $t,\mu,g$ are constants, and $a_{i}^{\dagger}$ (respectively, $a_{i}$) is the fermionic creation (respectively, annihilation) operator on site $i$. The results are shown in Fig. 1 for $n=40$ (main plot) and larger values of $n$ (inset). Here and throughout the paper, we use the sorted insertion algorithm [Crawford_Straaten_Wang_Parks_Campbell_Brierley_2021] to group terms into $k$-commuting sets. This algorithm sorts Hamiltonian terms by coefficient magnitude and then inserts them into $k$-commuting sets in a greedy manner (i.e., terms are inserted into the first mutually $k$-commuting group available, or inserted into a new group if no such group exists). This results in a number of groups which is plotted in blue in Fig. 1. As can be seen, the number of groups generally decreases as $k$ increases, although the behavior is not monotonic.
As discussed in  [Crawford_Straaten_Wang_Parks_Campbell_Brierley_2021], the number of groups is not the most suitable metric for evaluating the cost of measuring a Hamiltonian because terms with larger coefficients require more samples to measure. We borrow the example from [Crawford_Straaten_Wang_Parks_Campbell_Brierley_2021] to illustrate this: letting $H=4X_{1}+4X_{2}+Z_{2}+Z_{1}X_{2}$, the grouping $\left\{\{4X_{1},Z_{2}\},\{4X_{2},Z_{1}X_{2}\}\right\}$ results in the smallest number of groups, but the grouping $\left\{\{4X_{1},4X_{2}\},\{Z_{2}\},\{Z_{1}X_{2}\}\right\}$ requires fewer measurements because the terms with large coefficients are in the same group. (For an additional, non-trivial example, see Appendix B.) To quantify this, the authors of [Crawford_Straaten_Wang_Parks_Campbell_Brierley_2021] introduce the metric
@@
Here we have arranged the terms into $N$ groups with $m_{1}$, …., $m_{N}$ Paulis, and $c_{ij}$ is the coefficient of the $j$th Pauli in group $i$. This $\hat{R}$ value has been empirically shown to give a good approximation to the ratio $R:=M_{u}/M_{g}$ where $M_{u}$ is the number of shots needed to estimate an expectation value to precision $\epsilon$ without grouping the Pauli strings, and $M_{g}$ is the number of shots to estimate the expectation value to precision $\epsilon$ when the Pauli strings are grouped according to some grouping procedure $g$ [Crawford_Straaten_Wang_Parks_Campbell_Brierley_2021]. (Note that $\epsilon$ does not appear in $\hat{R}$ because it is a ratio.) The larger $\hat{R}$ is, the larger the reduction in measurement cost. For the remaining numerical experiments we adopt the $\hat{R}$ metric (6) to evaluate measurement reduction, but still show the number of groups for context.
Returning to the Fermi-Hubbard experiment in Fig. 1, we see that the $\hat{R}$ values shown in blue generally increase as $k$ increases, corresponding to a reduction in shot counts. Compared to $k=1$ (qubit-wise commuting), the measurement cost is cut in half for just $k=3$ and reduced by more than a factor of ten for $k=14$. Furthermore, we see the interesting property that $\hat{R}(k)$ attains its maximum at some $k^{*}<n$. In terms of practical experiments, this means that one can measure the energy of this Hamiltonian at the same cost as full commutativity using a diagonalization circuit with fewer gates than the one obtained via full commutativity. To further explore this, we repeat the experiment to find $k^{*}(n)$ for $n=80$ and $n=120$ qubits, and show the results in the inset of Fig. 1. Here, we see $k^{*}$ scales linearly with $n$, so while there are practical savings in measuring the energy the asymptotics are the same. In Sec. IV we explore this point further and discuss the asymptotics of $k^{*}(n)$ for several families of $n$-qubit Hamiltonians.
Using the same grouping algorithm and $\hat{R}$ metric, we test measurement reduction from $k$-commutativity for molecular Hamiltonians from HamLib [hamlib] listed in Table 1. Each molecule in HamLib includes many Hamiltonian instances, with qubit counts increasing past 100 qubits (i.e. many choices of activate space size are available). For this work we choose Hamiltonians of 10 to 16 qubits, testing 11 different molecules from the “standard” subdataset of the electronic structure dataset in HamLib. These molecules are all main group diatomics, where the ccPVDZ basis set and Bravyi-Kitaev encoding were used. The results are shown in Fig. 2. Here, we see a consistent increase in $\hat{R}$ as $k$ increases, demonstrating a reduction in cost for measuring expectation values. The reduction in the number of groups is also shown for context. We remark that the measurement reduction obtained by $k$-commutativity is similar using the Jordan-Wigner encoding instead of the Bravyi-Kitaev encoding — results are shown in Appendix C.
In practical applications, the choice of optimal $k$ depends largely on the particular quantum computer and overall measurement budget. If the computer is very noisy and one can afford running many circuits, it is best to choose a smaller $k$ to limit the depth of the diagonalizing circuit. On the other hand if the computer can implement operations with a high fidelity so that a larger diagonalizing circuit could be executed, it is best to choose a larger $k$ so that the overall measurement complexity is reduced.
@@
Table 1: Properties of molecular Hamiltonians obtained from HamLib [hamlib] used to test measurement reduction in $k$-commutativity (results shown in Fig. 2).
Figure 2: Results of numerical experiments computing the measurement cost $\hat{R}$ (6) (bottom panel) and number of groups (top panel) for molecular Hamiltonians listed in Table 1. For each molecule we evaluate the measurement cost of $k$-commutativity for $k\in\{1,n/4,n/2,3n/4,n\}$. As can be seen $\hat{R}$ increases as $k$ increases for all molecules, with the amount varying for different molecules.

### v3

@@ I Introduction
Figure 1: (Top panel) Measuring an $n=40\times 40$ Bacon-Shor Hamiltonian (5) by grouping terms into $k$-commuting sets for $1\leq k\leq n$. The left axis (blue) shows $\hat{R}$ defined in (6) as a function of $k$ and the right axis (green) shows the number of groups as a function of $k$. For measuring expectation values, higher $\hat{R}$ is better, and for applications such as Trotterization, a lower number of groups is better. As can be seen, the plot shows a “threshold” value $k^{*}<n$ for which $\hat{R}$ first reaches its maximum value and the number of groups first reaches its minimum value. (Bottom panel) The threshold value $k^{*}$ as a function of different Hamiltonian sizes $n$. In Sec. IV, we prove that $k^{*}(n)=\sqrt{n}$ for the Bacon-Shor Hamiltonian, in addition to discussing $k^{*}(n)$ for other Hamiltonian families. Here, our numerical results up to $n=40\times 40$ match the theory.
@@
For simplicity we refer to this as “measuring HH.” Without loss of generality we take $H$ to be an $n$-qubit Hamiltonian written as a linear combination of Pauli strings
@@
Motivated by the desire to have short-depth circuits due to hardware constraints, the notion of qubit-wise commutativity was developed [qubit-wise-commuting] [^1] Prior to this reference, the idea was introduced under the name of “tensor product basis sets” [Kandala_Mezzacapo_Temme_Takita_Brink_Chow_Gambetta_2017], although the term qubit-wise commutativity is most commonly used.. Two Pauli strings $P=\bigotimes_{i=1}^{n}p_{i}$ and $Q=\bigotimes_{i=1}^{n}q_{i}$ are said to qubit-wise commute if $p_{i},q_{i}]=0$ for all $i\in[n]$. If $P$ and $Q$ qubit-wise commute, they can be measured together in a depth-one quantum circuit. The catch is that, given a generic Hamiltonian (2), not all Pauli strings will qubit-wise commute. Approximate optimization algorithms are employed to group Paulis into qubit-wise commuting sets such that each set can be measured with a single depth-one circuit. This provides a trade off, decreasing circuit depth at the cost of increasing the number of circuits needed to be run.
Thus there are two notions of commutativity used in measuring $n$-qubit Hamiltonians: qubit-wise commuting which results in a larger number of circuits of additional depth $d=1$, and fully commuting which results in a fewer number of circuits of additional depth $d=O(n^{2}/\log n)$. In this paper we introduce a notion of commutativity which interpolates between these two. The idea is simply to consider commutativity on blocks of size $k\in\mathbb{N}$, and we call this $k$-commutativity. For example, we say Paulis $XXYY$ and $ZZXX$ two-commute since $[XX,ZZ]=0$ (the first two-qubit block) and $[YY,XX]=0$ (the second two-qubit block). Qubit-wise commutativity is recovered by $k=1$ and (full) commutativity is recovered by $k=n$. The advantage, with respect to measuring Hamiltonians, is the potential to reduce measurement complexity by allowing for increased circuit depth at intermediate $1\leq k\leq n$ values. An example of this is shown in Fig. 1, in which $k$-commutativity is applied to measuring the Hamiltonian of the $n$-qubit Bacon-Shor code. As shown in this plot, and proved later on in Sec. IV, this family of Hamiltonians features a “threshold value” $k^{*}=O(\sqrt{n})$ at which the number of groups is minimized, and the value of $\hat{R}$ (a ratio of the advantage of grouping compared to not grouping, to be defined later) is globally maximized. This interesting behavior, which we discuss in detail for several Hamiltonian families, demonstrates how our notion of $k$-commutativity can optimally reduce circuit depths when measuring expectation values by exploiting fine-grained commutativity structure, a crucial task for both near-term and fault-tolerant quantum computing.
@@ II $k$-commutativity
Let $P=P_{0}\otimes\cdots\otimes P_{n-1}$ be an $n$-qubit Pauli string. Given integers $a<b$, let $P_{a:b}$ denote $P$ restricted to $a$ through b1b-1, i.e. $P_{a}\otimes\cdots\otimes P_{b-1}$. For convenience we consider $P_{a},...,P_{b-1}$ to be identity if $a<0$, and similarly $P_{n},...,P_{b-1}$ to be identity if b>n1b>n-1. We define $k$-commutativity as follows.
@@
Two $n$-qubit Pauli strings $P=P_{0}\otimes\cdots\otimes P_{n-1}$ and $Q=Q_{0}\otimes\cdots\otimes Q_{n-1}$ are said to $k$-commute for integer $1\leq k\leq n$ if and only if $P_{ik:(i+1)k},Q_{ik:(i+1)k}]=0$ for all $i=0,...,\lfloor n/k\rfloor-1$. If $P$ and $Q$ $k$-commute, we write $P,Q]_{k}=0$.
Note that this definition assumes an ordering of the operators and throughout the paper we consider the ordering to be specified when evaluating $k$-commutativity. (If $k$ does not divide $n$, we can trivially pad the Pauli strings with identity operators to make this true.)
As mentioned, we note that $k=1$ corresponds to qubit-wise commutativity [qubit-wise-commuting] and $k=n$ corresponds to full commutativity. Due to its definition, $k$-commutativity inherits the usual properties of commutativity. In particular, $k$-commutativity is reflexive ($P,Q]_{k}=0$ implies $Q,P]_{k}=0$) but not transitive (if $P,Q]_{k}=0$ and $Q,R]_{k}=0$, then $P,R]_{k}$ could be zero or nonzero).
@@
Our next properties concern the number of Pauli strings that $k$-commute with a given Pauli string. For this we show a result for a slightly more general notion of $k$-commutativity in which the block sizes can vary. In particular, given two $n$-qubit Pauli strings $P$ and $Q$, and $k_{1},...,k_{m}\in\mathbb{N}$ such that $k_{1}+\cdots+k_{m}=n$, we say that $P$ and $Q$ $(k_{1},...,k_{m})$-commute if and only if $P_{a_{j}:b_{j}},Q_{a_{j}:b_{j}}]=0$ for every $j=1,\dots,m$, where $b_{j}=\sum_{\ell=1}^{j}k_{j}$ and $a_{j}=b_{j}-k_{j}$.
@@
For every $j=1,\dots,m$, define $b_{j}=\sum_{\ell=1}^{j}k_{j}$, and $a_{j}=b_{j}-k_{j}$. If $Q$ is a $n$-qubit Pauli that $(k_{1},\dots,k_{m})$-commutes with $P$, then for every $j$, we have $P_{a_{j}:b_{j}},Q_{a_{j}:b_{j}}]=0$. By [sarkar2021sets, Lemma 3], the number of distinct choices (modulo phase factors) for $Q_{a_{j}:b_{j}}$ is $4^{k_{j}}/2$. This gives that the total number of distinct possible choices (modulo phase factors) for $Q$ is $\prod_{j=1}^{m}(4^{k_{j}}/2)=4^{n}/2^{m}$. ∎
@@
Last, we note that the procedure for constructing a Clifford circuit to measure a set of $k$-commuting Paulis is identical to the case for (full) commutation [Crawford_Straaten_Wang_Parks_Campbell_Brierley_2021, Gokhale_Angiuli_Ding_Gui_Tomesh_Suchara_Martonosi_Chong_2020], just with the diagonalization procedure performed in $\lfloor n/k\rfloor$ blocks of size $k$ instead of one block of size $n$. For example, the two Paulis $X_{1}X_{2}X_{3}X_{4}$ and $Z_{1}Z_{2}Z_{3}Z_{4}$ 22-commute, so the diagonalization procedure would be called on each block of two qubits (in parallel) — namely to diagonalize $X_{1}X_{2}$ and $Z_{1}Z_{2}$ (the first block), and to diagonalize $X_{3}X_{4}$ and $Z_{3}Z_{4}$ (the second block). The full diagonalization circuit is then a tensor product of the two diagonalization circuits on each block.
@@ III Measurement reduction
As an initial test we consider the Hamiltonian of the Bacon-Shor code [ShorSchemeForReducingPRA1995, Bacon_2003] on an $M\times N$ lattice
$H=\sum_{i=1}^{M}S^{x}_{i}+\sum_{j=1}^{N}S^{z}_{j}$ (5)
where $S^{x}_{i}$ and $S^{z}_{j}$ are the stabilizers
$S_{i}^{x}=\bigotimes_{k=0}^{M-1}X_{i,k}X_{i+1,k},$
$S_{j}^{z}=\bigotimes_{k=0}^{N-1}Z_{k,j}Z_{k,j+1}.$
The results are shown in Fig. 1 for a square lattice of size $M=N=40$. Here and throughout the paper, we use the sorted insertion algorithm [Crawford_Straaten_Wang_Parks_Campbell_Brierley_2021] to group terms into $k$-commuting sets. This algorithm sorts Hamiltonian terms by coefficient magnitude and then inserts them into $k$-commuting sets in a greedy manner (i.e., terms are inserted into the first mutually $k$-commuting group available, or inserted into a new group if no such group exists). This results in a number of groups which is plotted in green in Fig. 1. As can be seen, there are values of $k$ at regular intervals which minimize the total number of groups needed to measure the Hamiltonian.
As discussed in Ref. [Crawford_Straaten_Wang_Parks_Campbell_Brierley_2021], the number of groups is not the most suitable metric for evaluating the cost of measuring a Hamiltonian because terms with larger coefficients require more samples to measure. We borrow the example from [Crawford_Straaten_Wang_Parks_Campbell_Brierley_2021] to illustrate this: letting $H=4X_{1}+4X_{2}+Z_{2}+Z_{1}X_{2}$, the grouping $\left\{\{4X_{1},Z_{2}\},\{4X_{2},Z_{1}X_{2}\}\right\}$ results in the smallest number of groups, but the grouping $\left\{\{4X_{1},4X_{2}\},\{Z_{2}\},\{Z_{1}X_{2}\}\right\}$ requires fewer measurements because the terms with large coefficients are in the same group. (For an additional, non-trivial example, see Appendix B.) To quantify this, the authors of [Crawford_Straaten_Wang_Parks_Campbell_Brierley_2021] introduce the metric
@@
Here we have arranged the terms into $N$ groups with $m_{1}$, …., $m_{N}$ Paulis, and $c_{ij}$ is the coefficient of the $j$th Pauli in group $i$. This $\hat{R}$ value has been empirically shown to give a good approximation to the ratio $R:=M_{u}/M_{g}$ where $M_{u}$ is the number of shots needed to estimate an expectation value to precision $\epsilon$ without grouping the Pauli strings, and $M_{g}$ is the number of shots to estimate the expectation value to precision $\epsilon$ when the Pauli strings are grouped according to some grouping procedure $g$ [Crawford_Straaten_Wang_Parks_Campbell_Brierley_2021]. (Note that $\epsilon$ does not appear in $\hat{R}$ because it is a ratio.) The larger $\hat{R}$ is, the larger the reduction in measurement cost. Since we primarily focus on measurement reduction in this work, in the remaining numerical experiments we adopt the $\hat{R}$ metric (6) to evaluate measurement reduction. However, minimizing the number of groups is more important for different applications — e.g., in the application of time evolution, fewer groups is beneficial for Trotter formulas because this minimizes the overall depth due to diagonalizing different groups. For these reasons, we show both $\hat{R}$ and the number of groups in remaining numerical experiments.
Returning to the Bacon-Shor Hamiltonian in Fig. 1, we see that $\hat{R}(k)$ attains its maximum value at some $k^{*}<n$. In terms of practical experiments, this means that one can measure the energy of this Hamiltonian at the same cost as full commutativity using a diagonalization circuit with fewer gates than the one obtained via full commutativity. To further explore this, we repeat the experiment to find $k^{*}(n)$ for various Hamiltonian sizes $n$, and show the results in the bottom panel of Fig. 1. Here, we see $k^{*}$ scales as $\sqrt{n}$, the linear size of the lattice. In Sec. IV we prove this asymptotic scaling and discuss the asymptotics of $k^{*}(n)$ for several other families of $n$-qubit Hamiltonians.
@@
N2 (3) 32 3637
Table 1: Properties of molecular Hamiltonians obtained from HamLib [hamlib] used to test measurement reduction via $k$-commutativity. Note that N2(3) simulates the triple-bond dissociation studied in [ibmn2] and taken from [sqdgithub]. Measurement cost results are shown in Fig. 2 for each molecule shown here.
Figure 2: Results of numerical experiments computing the measurement cost $\hat{R}$ (6) (bottom panel) and number of groups (top panel) for molecular Hamiltonians listed in Table 1. For each molecule we evaluate the measurement cost of $k$-commutativity for $k\in\{1,n/4,n/2,3n/4,n\}$. As can be seen, $\hat{R}$ generally increases as $k$ increases for all molecules, with the amount varying for different molecules.
Using the same grouping algorithm and $\hat{R}$ metric, we test measurement reduction from $k$-commutativity for molecular Hamiltonians from HamLib [hamlib] listed in Table 1. Each molecule in HamLib includes many Hamiltonian instances, with qubit counts increasing past 100 qubits (i.e. many choices of activate space size are available). For an initial study we choose Hamiltonians of 10 to 16 qubits, testing 11 different molecules from the “standard” subdataset of the electronic structure dataset in HamLib. These molecules are all main group diatomics, where the ccPVDZ basis set and Bravyi-Kitaev encoding were used. The results are shown in Fig. 2. Here, we see as expected that $\hat{R}$ is minimal for $k=1$ and maximal for $k=n$. The behavior of $\hat{R}$ for intermediate values $1\leq k\leq n$ is more interesting. While the general trend is that $\hat{R}$ increases as $k$ increases, the behavior is not monotonic. For example, with N2 we see that $\hat{R}(n/2)>\hat{R}(3n/4)$. In other cases, $\hat{R}$ is essentially equal for different $k$. For example, with O2 we see that $\hat{R}(n/2)\approx\hat{R}(3n/4)$. In cases such as this, it is of course advantageous to choose the smallest $k$ to reduce the complexity of the diagonalization circuit while retaining approximately the same measurement cost $\hat{R}$. We remark that the behavior in the number of groups as a function of $k$, shown in the top panel of Fig. 2, is qualitatively similar to the preceding discussion.
As a second numerical study, we calculate $\hat{R}$ for linear chains of Hydrogen on $n$ qubits (meaning there are $n/2$ Hydrogen atoms in the chain), a simple yet common system used to study scaling (see, e.g. [huggins2021efficient]). The results for these numerics are shown in Fig. 3. Unlike more structured Hamiltonians, e.g. the Bacon-Shor Hamiltonian (Fig. 1), here we do not see a “threshold” value of $k^{*}$ at which $\hat{R}$ saturates. We do, however, still see the general trend that $\hat{R}$ increases as $k$ increases, exhibiting the utility of $k$-commutativity to provide a tradeoff in number of shots vs. circuit depth when measuring expectation values. In terms of asymptotic measurement reduction for molecular Hamiltonians, it is known (see Table I of [huggins2021efficient]) that qubit-wise grouping ($k=1$) leads to $O(n^{4})$ groups with $O(1)$ measurement circuit depth, and full commuting ($k=n$) leads to $O(n^{3})$ groups with $O(n^{2}/\log n$) measurement circuit depth. Thus, using $k$-commuting with intermediate values $1\leq k\leq n$ allows one to smoothly interpolate between these two complexities, from a constant factor to a linear reduction. (We remark that these statements hold for molecular Hamiltonians — in Sec. IV we discuss other Hamiltonian families with different asymptotics.) In terms of practical measurement reduction for molecular Hamiltonians, as can be seen, this can lead to relatively large reductions in measuring energies — e.g., for the $n=64$ qubit linear Hydrogen chain, we see approximately a $\hat{R}(k)=200$ factor reduction in measuring the Hamiltonian relative to no grouping at $k=n/2=32$.
In practical applications, the choice of optimal $k$ depends largely on the particular quantum computer and overall measurement budget. If the computer is very noisy and one can afford running many circuits, it is best to choose a smaller $k$ to limit the depth of the diagonalizing circuit. On the other hand if the computer can implement operations with a high fidelity so that a larger diagonalizing circuit could be executed, it is best to choose a larger $k$ so that the overall measurement complexity is reduced. Whether $k$-commutativity will result in a constant-factor reduction or larger reduction in measurement complexity will depend largely on the particular Hamiltonian being studied. Thus far, we have seen two cases arising from numerical studies: (i) constant factor reductions in molecular Hamiltonians, and (ii) quadratic reductions in the Bacon-Shor Hamiltonian 5. We now seek to understand this behavior more rigorously.
Figure 3: Results of numerical experiments computing the measurement cost $\hat{R}$ (6) for Hydrogen chains on $n$ qubits. Each molecule consists of a linear chain of $n/2$ Hydrogen atoms. For each we evaluate the measurement cost for $1\leq k\leq n$. As can be seen, increasing $k$ generally results in a higher $\hat{R}$, or equivalently lower measurement cost, although the behavior is not monotonic, and $k$-commutativity enables a relatively smooth interpolation between single-qubit commuting and (fully) commuting.